Find Pluto's

a) minimum distance from the sun

b) maximum distance from the sun

Note:

G= 6.67x10-11 Nm2/kg2

Conversion:




1
Expert's answer
2022-02-17T04:35:00-0500

a)

"q=a(1-e)=29.67~AU,"

b)

"Q=a(1+e)=49.31~AU."